# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ 4 Crossing densities of graphons
Let $W\in\mathcal{W}_{0}$. We shall exhibit a sequence of graphs $G_{1},G_{2},\dots$ such that $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ and $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W),\ \lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$. Suppose for a moment that such a sequence exists. Then, it must be the case that $\overline{\mathop{\mathrm{cd}}}$ and the functional $\overline{g}$ mentioned in Theorem 3.5 are one and the same, and this also true for $\mathop{\mathrm{cd}}$ and $g$. Since $\overline{g}$ and $g$ are continuous with respect to $d_{\square}$, so too are $\overline{\mathop{\mathrm{cd}}}$ and $\mathop{\mathrm{cd}}$. In order to construct this graph sequence, we need some preliminaries.
For any $W\in\mathcal{W}$ and any partition $\mathcal{P}=\{S_{1},S_{2},\dots,S_{n}\}$ of the unit interval into measurable sets, let $W_{\mathcal{P}}^{\text{diag}}$ be the graphon such that $W_{\mathcal{P}}^{\text{diag}}(x,y)=W_{\mathcal{P}}(x,y)$ if $x$ and $y$ lie in different parts of $\mathcal{P}$, and $W_{\mathcal{P}}^{\text{diag}}(x,y)=0$ otherwise. The advantage that we gain from using $W_{\mathcal{P}}^{\text{diag}}$ instead of $W_{\mathcal{P}}$ is that if $\mathcal{P}$ is a partition of $[0,1]$ into intervals of the same length, then $W_{\mathcal{P}}^{\text{diag}}=W_{G}$ for some graph $G$. Note that, if we let $\mathcal{P}_{n}$ be the partition into intervals of length $1/n$ for every positive integer $n$, then each pair of distinct points in $[0,1]$ is separated by all but a finite amount of the $P_{i}$’s. As a consequence, the sequence $W_{\mathcal{P}_{1}},W_{\mathcal{P}_{2}},\dots$ converges to $W$ almost everywhere. Observe that for every point $(x,y)\in[0,1]^{2}$ with $x\neq y$ there exists an integer $N$ such that $W_{\mathcal{P}_{n}}(x,y)=W_{\mathcal{P}_{n}}^{\text{diag}}(x,y)$ whenever $n\geq N$. This implies that the sequence $W_{\mathcal{P}_{1}}^{\text{diag}},W_{\mathcal{P}_{2}}^{\text{diag}},\dots$ converges to $W$ almost everywhere too. Let $G_{1},G_{2},\dots$ be edge weighted graphs such that $W_{G_{n}}=W_{\mathcal{P}_{n}}^{\text{diag}}$ for every $n$, then, $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ almost everywhere and Theorem 4.2 implies that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W)$ and $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$. Since convergence almost everywhere implies convergence with respect to $d_{\square}$, the result follows. ∎
@@
The crossing density and the rectilinear crossing density of graphons are invariant parameters.
We point out that this can also be deduced in a more natural way without having to recur, indirectly, to crossing densities of graphs.
@@
Clearly, if $c\geq 0$ and $W_{1}$ and $W_{2}$ are graphons with $W_{1}=cW_{2}$, then $\mathop{\mathrm{cd}}(W_{1})=c^{2}\cdot\mathop{\mathrm{cd}}(W_{2})$ and $\overline{\mathop{\mathrm{cd}}}(W_{1})=c^{2}\cdot\overline{\mathop{\mathrm{cd}}}(W_{2})$. Crossing numbers of random graphs have received considerable attention (see, for example, [39]). A sequence $G_{1},G_{2},\dots$ of edge weighted graphs is said to be quasi-random with density $p$ if the orders of the graphs go to infinity with $n$ and $W_{G_{1}},W_{G_{2}},\dots$ converges to $W^{p}$. There are many other equivalent (and often more practical) definitions of quasi-randomness (see [11, 29]), but we restrict ourselves to this one because it is the most convenient for us. The results in [18] directly imply that the rectilinear crossing densities of any quasi-random sequence of graphs converge. We can show that the same is true for the crossing density; in fact, if we write $\overline{K}=\overline{\mathop{\mathrm{cd}}}(W_{1})$ and $K=\mathop{\mathrm{cd}}(W_{1})$, then $\mathop{\mathrm{cd}}(W^{p})=Kp^{2}$ and $\overline{\mathop{\mathrm{cd}}}(W^{p})=\overline{K}p^{2}$, whence Theorem 4.4 yields the following.
@@
Let $p\in[0,1]$. Then, for any sequence $G_{1},G_{2},\dots$ of edge weighted graphs that is quasi-random with density $p$, we have that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(G_{n})=\overline{K}p^{2}$ and $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(G_{n})=Kp^{2}$.
@@
The sequence $W_{1},W_{2},\dots$ converges to $W_{\mathcal{P}}$ almost everywhere by the choice of $\phi$, hence $\overline{\mathop{\mathrm{cd}}}(W_{\mathcal{P}})=\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\geq\overline{\mathop{\mathrm{cd}}}(W)$. The result for crossing densities follows by an analogous argument. ∎
@@
It might seem inconvenient that the definitions of $\mathcal{F}$ and $\mathcal{D}_{\mathcal{W}}$ allow for "degenerate" configurations. However, it is possible to add several more restrictions to these families without altering the crossing densities. Below, an example of this is provided.
@@
Combining the ideas in the proofs of Theorems 4.3 and 4.4, we can obtain a sequence $f_{1},f_{2},\dots$ of elements of $\mathcal{F}$, each of which arises from a rectilinear drawing of a graph (as did $f_{\mathcal{D}}$ in our proof that $\overline{\mathop{\mathrm{cd}}}(W_{G})\leq\overline{\mathop{\mathrm{cd}}}+1/n$ for every $n$ vertex graph $G$), and such that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W,f_{n})=\overline{\mathop{\mathrm{cd}}}(W)$. For every maximal subinterval of $[0,1]$ that is mapped to a single point by $f_{n}$, we can reconstruct $f_{n}$ so that its restriction to the subinterval is instead measure preserving and its image is an arbitrarily small disk. This modification can be carried out so that the rectilinear crossing density increases by an arbitrarily small amount. Repeating this process for every subinterval with the above property and then for every $n$, we arrive at a new sequence $f_{1}^{\prime},f_{2}^{\prime},\dots$ of nice elements of $\mathcal{F}$ such that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W,f_{n}^{\prime})=\overline{\mathop{\mathrm{cd}}}(W)$. ∎
We mentioned in the introduction that the crossing number of complete graphs is connected to Sylvester’s four point problem. To finish this section, we describe how this connection is, in some sense, a particular case of the interplay between the rectilinear crossing density of graphons and the rectilinear crossing number of graphs.
Given a bounded region $R\subset\mathbb{R}^{2}$ of positive measure, let $\lambda_{R}$ be the probability measure that arises from restricting the Lebesgue measure to $R$ and then normalizing. For any probability measure $P$ on the Lebesgue $\sigma$-algebra of $\mathbb{R}^{2}$, let $\text{c}(P)$ denote the probability that four points sampled independently from $P$ can be labeled by $a,b,c,d$ so that $I(a,b,c,d)=1$ (recall that $I$ was defined at the beginning of this section). It was shown by Scheinerman and Wilf [38] (although it is worded in a slightly different manner in their paper) that
@@
Now, for $f\in\mathcal{F}$, let $\lambda_{f}$ denote the corresponding pushforward probability measure on $\mathbb{R}^{d}$ (i.e., $\lambda_{f}(A)=\lambda(f^{-1}(A))$ for every measurable $A\subseteq\mathbb{R}^{2}$). Every probability measure of the form $\lambda_{R}$ is standard, and thus can be written as $\lambda_{f}$ for some nice $f\in\mathcal{F}$. Furthermore, if $f$ is nice then $\lambda_{f}=\lambda_{R}$ for some $R$. This implies that
@@
where $f$ ranges over all nice elements of $\mathcal{F}$. On top of this, it is not hard to see that $24\overline{\mathop{\mathrm{cd}}}(W^{1},f)=\text{c}(\lambda_{f})$ as long as the set $f^{-1}(\ell)\subset[0,1]$ has measure 00 for every line $\ell$ on the plane. This observation, along with by Theorem 4.11, implies that
@@
5 Further research
Unlike the algorithm from [18], our drawing algorithm from Theorem 1.2 is not deterministic. We leave the following open problem:
@@ 5 Further research
While Theorem 4.4 shows that the graphon parameters $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ are indeed the "correct" discrete analogs of the rectilinear crossing density and the crossing density of graphs, there are several ways in which we could have defined them. Indeed, as we exemplified at the end of the previous section, minor tweaks to the definitions of $\mathcal{F}$ and $\mathcal{D}_{\mathcal{W}}$ will often have no influence whatsoever on the graphon parameters $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ (see Theorem 4.11). Since the definition of $\mathcal{D}_{\mathcal{W}}$ is rather artificial (it is made up of all the simple graphon drawings such that $I_{\mathcal{D}}$ measurable), it makes sense to ask whether this family can be substituted for a more reasonable set of functions from $[0,1]^{2}$ to $\mathcal{C}$ without altering $\mathop{\mathrm{cd}}$.
@@
For any $p\in[0,1]$, the set of graphons with $||W||_{1}=p$ is clearly closed with respect to $d_{\square}$. Since $(\widetilde{\mathcal{W}_{0}},\delta_{\square})$ is compact and the crossing densities are invariant, the following problem makes sense.
Problem 5.2.
For every $p\in[0,1]$, determine two graphon $W_{1}$ and $W_{2}$ with $||W_{1}||_{1}=||W_{2}||_{1}=p$ such that
@@
Corollary 4.10 provides a solution to the variant of this problem where the minimums are substituted by maximums. We expect both of the above questions to be difficult to answer.

### v4

@@ 4.2 Continuity with respect to $d_{\square}$ and connection to graphs
Let $W\in\mathcal{W}_{0}$. We shall exhibit a sequence of graphs $G_{1},G_{2},\dots$ such that $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ and $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W),\ \lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$. Suppose for a moment that such a sequence exists. Then, it must be the case that $\overline{\mathop{\mathrm{cd}}}$ and the functional $\overline{g}$ mentioned in Theorem 3.5 are one and the same, and this also true for $\mathop{\mathrm{cd}}$ and $g$. Since $\overline{g}$ and $g$ are continuous with respect to $d_{\square}$, so too are $\overline{\mathop{\mathrm{cd}}}$ and $\mathop{\mathrm{cd}}$.
In order to construct this graph sequence, we introduce some notation. For any $W\in\mathcal{W}$ and any partition $\mathcal{P}=\{S_{1},S_{2},\dots,S_{n}\}$ of the unit interval into measurable sets, let $W_{\mathcal{P}}^{\text{diag}}$ be the graphon such that $W_{\mathcal{P}}^{\text{diag}}(x,y)=W_{\mathcal{P}}(x,y)$ if $x$ and $y$ lie in different parts of $\mathcal{P}$, and $W_{\mathcal{P}}^{\text{diag}}(x,y)=0$ otherwise. The advantage that we gain from using $W_{\mathcal{P}}^{\text{diag}}$ instead of $W_{\mathcal{P}}$ is that if $\mathcal{P}$ is a partition of $[0,1]$ into intervals of the same length, then $W_{\mathcal{P}}^{\text{diag}}=W_{G}$ for some graph $G$. Note that if we let $\mathcal{P}_{n}$ be the partition into intervals of length $1/n$ for every positive integer $n$, then each pair of distinct points in $[0,1]$ is separated by all but a finite amount of the $P_{i}$’s. As a consequence, the sequence $W_{\mathcal{P}_{1}},W_{\mathcal{P}_{2}},\dots$ converges to $W$ almost everywhere. Observe that for every point $(x,y)\in[0,1]^{2}$ with $x\neq y$ there exists an integer $N$ such that $W_{\mathcal{P}_{n}}(x,y)=W_{\mathcal{P}_{n}}^{\text{diag}}(x,y)$ whenever $n\geq N$. This implies that the sequence $W_{\mathcal{P}_{1}}^{\text{diag}},W_{\mathcal{P}_{2}}^{\text{diag}},\dots$ converges to $W$ almost everywhere too. Let $G_{1},G_{2},\dots$ be edge-weighted graphs such that $W_{G_{n}}=W_{\mathcal{P}_{n}}^{\text{diag}}$ for every $n$. Then, $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ almost everywhere and Lemma 4.2 implies that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})=\overline{\mathop{\mathrm{cd}}}(W)$ and $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(W_{n})=\mathop{\mathrm{cd}}(W)$. Since convergence almost everywhere implies convergence with respect to $d_{\square}$, the result follows. ∎
@@
The crossing density and the rectilinear crossing density of graphons are invariant.
We point out that this corollary can also be deduced in a more direct manner without having to recur to crossing densities of graphs and Theorem 3.5.
4.3 Some other properties of the crossing densities
@@
Clearly, if $c\geq 0$ and $W_{1}$ and $W_{2}$ are graphons with $W_{1}=cW_{2}$, then $\mathop{\mathrm{cd}}(W_{1})=c^{2}\cdot\mathop{\mathrm{cd}}(W_{2})$ and $\overline{\mathop{\mathrm{cd}}}(W_{1})=c^{2}\cdot\overline{\mathop{\mathrm{cd}}}(W_{2})$. Crossing numbers of random graphs have received considerable attention (see, for example, [47]). A sequence $G_{1},G_{2},\dots$ of edge-weighted graphs is said to be quasi-random with density $p$ if the orders of the graphs go to infinity with $n$ and $W_{G_{1}},W_{G_{2}},\dots$ converges to $W^{p}$. There are many other equivalent (and often more practical) definitions of quasi-randomness (see [12, 34]), but this one is the best suited for our purposes. The results in [21] directly imply that the rectilinear crossing densities of any quasi-random sequence of graphs converge. We can show that the same is true for the crossing density. In fact, if we write $\overline{K}=\overline{\mathop{\mathrm{cd}}}(W^{1})$ and $K=\mathop{\mathrm{cd}}(W^{1})$, then $\mathop{\mathrm{cd}}(W^{p})=Kp^{2}$ and $\overline{\mathop{\mathrm{cd}}}(W^{p})=\overline{K}p^{2}$, and so Theorem 4.4 yields the following.
@@
Let $p\in[0,1]$. Then, for any sequence $G_{1},G_{2},\dots$ of edge-weighted graphs that is quasi-random with density $p$, we have that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(G_{n})=\overline{K}p^{2}$ and $\lim_{n\rightarrow\infty}\mathop{\mathrm{cd}}(G_{n})=Kp^{2}$.
@@
The sequence $W_{1},W_{2},\dots$ converges to $W_{\mathcal{P}}$ almost everywhere by the choice of $\phi$, hence $\overline{\mathop{\mathrm{cd}}}(W_{\mathcal{P}})=\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W_{n})\geq\overline{\mathop{\mathrm{cd}}}(W)$. The result for crossing densities follows from an analogous argument. ∎
@@
It might seem inconvenient that the definitions of $\mathcal{F}$ and $\mathcal{D}_{\mathcal{W}}$ allow for "degenerate" configurations. However, it is possible to add several more restrictions to these families without altering the crossing densities. We provide an example of this.
@@
Combining the ideas in the proofs of theorems 4.3 and 4.4, we can obtain a sequence $f_{1},f_{2},\dots$ of elements of $\mathcal{F}$, each of which arises from a rectilinear drawing of a graph (as did $f_{\mathcal{D}}$ in our proof that $\overline{\mathop{\mathrm{cd}}}(W_{G})\leq\overline{\mathop{\mathrm{cd}}}(G)+1/n$ for every $n$ vertex graph $G$), and such that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W,f_{n})=\overline{\mathop{\mathrm{cd}}}(W)$. Each of these elements of $F$ maps each of the intervals in some partition of $[0,1]$ to a distinct point on the plane. We can rescale $f_{n}$ so that the points corresponding to the images of any two of these intervals are arbitrarily far away from each other. For every maximal subinterval of $[0,1]$ that is mapped to a single point by $f_{n}$, we can now reconstruct $f_{n}$ so that its restriction to the subinterval is instead measure preserving and its image is a disk. This modification can be carried out so that the rectilinear crossing density increases by an arbitrarily small amount. Repeating this process for every subinterval with the above property and then for every $n$, we arrive at a new sequence $f_{1}^{\prime},f_{2}^{\prime},\dots$ of nice elements of $\mathcal{F}$ such that $\lim_{n\rightarrow\infty}\overline{\mathop{\mathrm{cd}}}(W,f_{n}^{\prime})=\overline{\mathop{\mathrm{cd}}}(W)$. ∎
We mentioned in the introduction that the crossing number of complete graphs is connected to Sylvester’s four point problem. To finish this section, we describe how this connection is, in some sense, a consequence of the interplay between rectilinear crossing densities of graphons and rectilinear crossing numbers of graphs.
Given a bounded region $R\subset\mathbb{R}^{2}$ of positive measure, let $\lambda_{R}$ be the probability measure that arises from restricting the Lebesgue measure to $R$ and then normalizing. For any probability measure $P$ on the Lebesgue $\sigma$-algebra of $\mathbb{R}^{2}$, let $\text{c}(P)$ denote the probability that four points sampled independently from $P$ can be labeled by $a,b,c,d$ so that $I(a,b,c,d)=1$ (that is, $\text{conv}\{a,b\}\cap\text{conv}\{c,d\}\neq\emptyset$). It was shown by Scheinerman and Wilf [46] (although it is worded in a slightly different manner in their paper) that
@@
Now, for $f\in\mathcal{F}$, let $\lambda_{f}$ denote the corresponding pushforward probability measure on $\mathbb{R}^{d}$ (i.e., $\lambda_{f}(A)=\lambda(f^{-1}(A))$ for every measurable $A\subseteq\mathbb{R}^{2}$). Every probability measure of the form $\lambda_{R}$ corresponds to a standard probability space, and thus each $\lambda_{R}$ can be written as $\lambda_{f}$ for some nice $f\in\mathcal{F}$. Furthermore, if $f$ is nice then $\lambda_{f}=\lambda_{R}$ for some $R$. This implies that
@@
where $f$ ranges over all nice elements of $\mathcal{F}$. On top of this, it is not hard to see that $24\overline{\mathop{\mathrm{cd}}}(W^{1},f)=\text{c}(\lambda_{f})$ for all nice $f$ (since the set $f^{-1}(\ell)\subset[0,1]$ has measure 00 for every line $\ell$ on the plane). This observation, along with Theorem 4.11, implies that
@@
5 Open problems
  • Unlike the algorithm from [21], our drawing algorithm in Theorem 1.2 is not deterministic.
@@ 5 Open problems
  • While Theorem 4.4 shows that the graphon parameters $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ are indeed the "correct" discrete analogs of the rectilinear crossing density and the crossing density of graphs, there are several ways in which we could have defined them. Indeed, as we exemplified at the end of the previous section, minor tweaks to the definitions of $\mathcal{F}$ and $\mathcal{D}_{\mathcal{W}}$ will often have no influence whatsoever on the graphon parameters $\mathop{\mathrm{cd}}$ and $\overline{\mathop{\mathrm{cd}}}$ (see Theorem 4.11). Since the definition of $\mathcal{D}_{\mathcal{W}}$ is rather artificial (the set is made up of all the simple graphon drawings such that $I_{\mathcal{D}}$ measurable), it makes sense to ask whether this family can be substituted by a more natural set of functions from $[0,1]^{2}$ to $\mathcal{C}$ without altering $\mathop{\mathrm{cd}}$.
  • As mentioned near the end of Section 4.1, we do not know whether for every graphon $W$ there is an $f\in\mathcal{F}$ such that $\overline{\mathop{\mathrm{cr}}}(W)=\overline{\mathop{\mathrm{cr}}}(W,f)$ or a $\mathcal{D}\in\mathcal{D}_{\mathcal{W}}$ such that $\mathop{\mathrm{cd}}(W)=\mathop{\mathrm{cd}}(W,\mathcal{D})$. We believe that the answer to the first of these questions is negative, and we make the following conjecture.
    Conjecture 5.2.
    There exists no $f\in\mathcal{F}$ such that $\overline{\mathop{\mathrm{cd}}}(W^{1},\mathcal{D})=\overline{\mathop{\mathrm{cd}}}(W^{1},f)$. In particular, the infimum in
    $\lim_{n\rightarrow\infty}\frac{\overline{\mathop{\mathrm{cr}}}(K_{n})}{\binom{n}{4}}=\inf_{R}c(\lambda_{R})$
    cannot be substituted by a minimum.
@@
  • For any $p\in[0,1]$, the set of graphons with $||W||_{1}=p$ is clearly closed with respect to $d_{\square}$. Since $(\widetilde{\mathcal{W}_{0}},\delta_{\square})$ is compact and the crossing densities are invariant, the following problem is well posed.
    Problem 5.3.
    For every $p\in[0,1]$, find two graphon $W_{1}$ and $W_{2}$ with $||W_{1}||_{1}=||W_{2}||_{1}=p$ such that
@@
    Corollary 4.10 provides an answer to the variant of this problem where the minimums are substituted by maximums. We expect both of the above questions to be difficult to answer, but we believe that doing so might be an important step towards determining the so-called midrange crossing constant [^12] Let $\kappa(n,e)$ denote the minimum crossing number among all $n$-vertex graphs $G$ with $e$ edges. It was shown by Pach, Spencer and Tóth [39] that there exists a constant $\kappa$ (the midrange crossing constant) such that, under the assumption that $n\ll e\ll n^{2}$, $\lim_{n\rightarrow\infty}\kappa(n^{2}/e^{3})=\kappa$. An analogous result holds for rectilinear crossing numbers..
    Let $\mathbb{S}^{2}$ denote a $2$-dimensional sphere in $\mathbb{R}^{3}$ with total surface area $1$. We will now describe a graphon defined on $\mathbb{S}^{2}$ instead of $[0,1]$, but we remark that it can be transformed to an equivalent graphon in $\mathcal{W}_{0}$ by means of a measure preserving transformation, and so the theory we have developed still applies to it. Fix $p\in[0,1]$ and let $\tau_{p}$ denote the unique real number such that, for any $x\in\mathbb{S}^{2}$, the intersection of $\mathbb{S}^{2}$ with the ball of radius $\tau_{p}$ centered at $x$ has surface area $p$. Define $W_{\mathbb{S}^{2},p}:\mathbb{S}^{2}\times\mathbb{S}^{2}\rightarrow\mathbb{[}0,1]$ by setting $W_{\mathbb{S}^{2},p}(x,y)=1$ if $x$ and $y$ are at distance at most $\tau_{p}$ and $W_{\mathbb{S}^{2},p}(x,y)=0$ otherwise. Inspired by the main result of [16], we make the following conjecture.
    Conjecture 5.4.
    For every $p\in[0,1]$, we have that
    $\operatorname{cd}(W_{\mathbb{S}^{2},p})=\min_{W\in\mathcal{W}_{0},||W||_{1}=p}\mathop{\mathrm{cd}}(W).$
  • Finally, motivated by the Kuratowski-Wagner Theorem [30, 51] (which states that a graph is planar if and only if it doesn’t contain a subdivision of $K_{5}$ or $K_{3,3}$ as a subgraph), we ask the following question.
    Problem 5.5.
    Let $W$ be a graphon. What can be said about $\operatorname{cd}(W)$ and $\overline{\operatorname{cd}}(W)$ just from knowing the homomorphism densities $t(F,W)$ where $F$ ranges over all subdivisions of $K_{5}$ and $K_{3,3}$?